(I) A square matrix with the characteristic polynomial 14 – 413 +212 – +3 is invertible. [ 23] (II) Matrix in Z5 has two distTrue or false. Please justify why true or why false also
Answer #1

polynomial A į (-A+4A-ZA+I) a square matrix satisfied its characteristic 44- 41²+2 1 + 3 = Hene A-HA? UZA- A+B=0 (AC- & A²+2FALSE similar matrix have not the same corresponding eigenvalue. eigen spare For und are The vector 1.:) los similar with eig